# PromptRPA: Generating Robotic Process Automation on Smartphones from Textual Prompts

### v1

@@ 2. Clear cache
QQ Message 2 1. Share QQ screen
@@ 2. Provide feedback on an issue
TikTok Social Media 3 1. Cancel the monthly payment
@@ 2. View TikTok shop
Alipay Pay 4 1. Set the notification sound
@@ 2. View Alipay points
Weibo Social Media 5 1. Check creation earnings
@@ 2. View corrected followers
Beautiful Weather Weather 366 1. Update the weather every 3 hours
2. Observe the weather in Beijing
Safe and Sound Lifestyle 392 1. Set the shipping address
2. Enable the night mode do not disturb
Snail Sleep Lifestyle 418 1. Set the goal as stress relief
2. View one’s own sleep talk
Film Encyclopedia Entertainment 671 1. Join the user experience program
2. Bind QQ with the Film Encyclopedia
HUAWEI System Tools System - 1. Turn off Pure Mode
@@ 5.2. Experiment Method
In our experiment, we evaluated three configurations of our system:
  • Standard PromptRPA.
  • PromptRPA(AK): PromptRPA enhanced with accumulated knowledge.
  • PromptRPA(AK+UI): PromptRPA further enhanced with both accumulated knowledge and user intervention.
The experiment assessed PromptRPA’s performance at individual stages and across the entire pipeline. Notably, in a single stage where a task was not fully completed, without user intervention, the knowledge that PromptRPA could accumulate was quite limited. Therefore, PromptRPA(AK) was evaluated solely on its entire pipeline performance. When evaluating the performance of individual stages, we used ground-truth inputs, ensuring that any output deviation stemmed exclusively from the current stage, thus eliminating errors from preceding stages. Specifically, in the instruction generation stage, we input 100 ground-truth function and step descriptions; for the operation mapping stage, we input 100 ground-truth instruction lists. In evaluating the information collection stage and the entire pipeline, the input comprised 2,500 original textual prompts. To minimize the impact of input order on knowledge accumulation, we randomly shuffled the inputs and conducted five rounds of testing to calculate the average performance.
We selected appropriate baselines for comparison at both individual stages and the entire process. For the information collection stage, manual queries on Google, Bing, and Baidu served as the baseline. For instruction generation, we used Seq2act (Li et al., 2020a) as the baseline. For the operation mapping stage, we employed strict text matching as the baseline. In assessing the entire pipeline, due to the lack of existing work supporting open-domain mobile tasks, our baseline was instructions generated purely by GPT-4, followed by strict text matching for operation mapping.
@@ 5.3. Experiment Results
The experiment results are organized based on the three stages within PromptRPA as well as the performance of the entire pipeline.
@@ 5.3.1. Information Collection
This stage focuses on evaluating whether the system can successfully generate a function description and step description based on a user-provided textual prompt.
Among the 2,500 textual prompts, 100 contained complete step information and required no further online retrieval. Users accepted the analyzed results for 97.00% of these prompts, with a few exceptions attributed to the LLM’s data fabrication issues. The remaining 2,400 prompts were used to assess the top 5 tutorials retrieved by the system. The criteria for a ‘hit’ was whether users deemed a tutorial matching the ground-truth function description, regardless of on-phone executability. A tutorial at the top indicated that the system could automatically select it, eliminating the need for user intervention.
@@
Table 3. Performance of various search engines in retrieving tutorials with ground-truth function descriptions of RPA task models.
@@
Table 4. Performance of information collection.
@@
Manual(Baidu) 97.21 1.42 (sd=2.76) 87.96 -
PromptRPA 98.83 1.27 (sd=0.91) 89.38 88.38
PromptRPA(AK+UI) 100.00 1.22 (sd=0.74) 90.83 99.21
However, prompts provided by users often lack complete accuracy, leading to a decrease in effectiveness when using the original prompts for manual searches on these websites. Using them as baselines, we tested the performance of PromptRPA. As demonstrated in Table 4, our system improved the Top-5 hit rate from 97.21% (Baidu) to 100%, while achieving a Top-1 hit rate of 90.83%. On average, each prompt required 1.15 user interventions (sd = 0.52), which is the number of conversations between the user and PromptRPA. Remarkably, without user interventions, our system demonstrated a Top-1 hit rate of 89.38%, surpassing the baseline performance.
After the agent matched function descriptions with online tutorials, it filtered concise step descriptions from complex web data. The accuracy of step descriptions refers to the correct extraction from tutorial web pages, ensuring they are free from fabrication, omissions, or errors. When the system automatically selected the first candidate, the accuracy was 88.38%. When users could select from five candidates, the accuracy rose to 99.21%.
@@ 5.3.2. Instruction Generation
Partial Accuracy(%) Complete Accuracy(%)
Seq2act (Li et al., 2020a) 70.13 54.00
PromptRPA 84.55 73.00
PromptRPA(AK+UI) 100.00 100.00
This stage evaluates the system’s ability to transform step descriptions into a list of instructions in the predefined format.
Using the ground-truth step descriptions of the selected 100 RPA task models as input, we evaluated correctness using two metrics, focusing solely on textual interpretation:
  • Complete Accuracy: The percentage of RPA task models where all generated instructions match the ground truth.
We used the Seq2Act model (Li et al., 2020a), a transformer trained using the ANDROIDHOWTO dataset, as our baseline (after translating step descriptions into English). Table 5 reveals that PromptRPA achieved 100% accuracy for both complete and partial metrics. On average, achieving perfection required 0.93 user interventions per RPA task (sd=1.34). It’s worth noting that while perfection necessitated user interventions, the subsequent pipeline stages are designed to accommodate minor errors, reducing the need for strict user oversight. Without user intervention, the system attained 84.55% partial and 73.00% complete accuracy. In cold-start scenarios, PromptRPA has achieved performance comparable to that achieved by related works (Li et al., 2020a; Vu et al., 2023), which require substantial training data (at least 10K-level).
@@ 5.3.3. Operation Mapping
Table 6. Performance of operation mapping.
Prediction Accuracy(%) Success Rate(%)
Text only 61.49 44.00
PromptRPA 81.66 76.00
PromptRPA(AK+UI) 100.00 96.00
This stage evaluates the system’s ability to execute operations on smartphones, guided by instructions. We used the ground-truth instruction lists of the 100 RPA task models for testing.
Our baseline for comparison utilized strict text matching to identify operation widgets. Table 6 shows the results. The Grounding Agent’s prediction accuracy is measured by the ratio of operations that align with the ground-truth operation sequences. With user intervention (i.e., users can demonstrate any operations), the success rate reached 96.00%. The 4 unsuccessful RPA tasks were infeasible in the test environment. The Assessment Agent shows a 77.27% precision and a 97.06% recall rate for identifying instances requiring user intervention. On average, each RPA task required 0.68 user interventions (sd = 0.85). Even without user assistance, the success rate was 76.00%, far surpassing the baseline.
Notably, not all instructions were executable on smartphones, usually due to obsolescence from application updates. Our tests revealed that 81.33% of the ground-truth instructions were feasible in the test environment. Thus, our system does more than merely interpret the original instructions; it also performs its own assessments and predictions based on encountered mobile pages.
@@ 5.3.4. The Entire Pipeline
Table 7. Success rate (%) across various stages of PromptRPA.
@@
PromptRPA 88.60 65.04 72.80
PromptRPA(AK) 92.72 72.44 83.68
PromptRPA(AK+UI) 99.12 99.12 95.21
To evaluate the overall effectiveness of our system, we input 2,500 textual prompts into the pipeline, which comprises information collection, instruction generation, and operation mapping. We did not use UGIF (Venkatesh et al., 2022) as a baseline, because our tasks were not included in its tutorial collection. We did not use Voicify (Vu et al., 2023) either, because its direct invocation requires the apps’ code to contain deep links, but most of our tasks did not satisfy it. Instead, we established our baseline using GPT-4 for direct instruction generation from user inputs, coupled with strict text matching for operations.
Are shown in Table 7, the success rate for each stage is strictly determined by the complete accuracy achieved up to that stage. Importantly, the overall success rate is not a simple product of success rates at each stage, due to the system’s built-in flexibility for error correction in downstream stages. For instance, the operation mapping stage doesn’t strictly follow the generated instructions; it also utilizes the agent’s explorative reasoning. With user intervention, the average number of interventions per task is 1.66 (sd = 1.18). Notably, this number is smaller than the sum of interventions required for each stage as reported earlier, which shows that the accumulation of knowledge from completed tasks can reduce user interventions.
@@ 5.4. Error Analysis
We analyzed the errors that occurred in the three stages of PromptRPA.
@@ 5.4.1. Information Collection
Our system effectively matches tutorial web pages but falls short in content extraction. Two main issues affect the performance. First, web pages containing excessively long text compromise the LLM’s extraction capabilities, leading to information loss and increased risk of generating inaccurate content. Second, the LLM also tends to generate fabricated step descriptions when interpreting the textual prompts.
To address these challenges, we have considered two options for improving text quality. Initially, we considered implementing a specialized assessment agent, similar to the one used in the operation mapping, to manage data quality. However, we chose a human-in-the-loop approach, considering the efficiency of manual text assessment and the extra time needed for an agent-based solution. Our current system allows users to manually select high-quality step descriptions for subsequent stages.
@@ 5.4.2. Instruction Generation
Figure 11. Errors of automatic instruction generation.
Figure 12. Errors of automatic operation mapping.
Our system’s automatic instruction generation achieves a partial accuracy of 84.55%, with errors distributed across six main categories as shown in Figure 12. These errors are generally resolved through simple user corrections. Specifically, 43.75% of errors are due to instructions that go beyond the existing definition (Table 4.1.3). Within this subset, 27.08% involve the LLM using near-synonyms (e.g., “drag down” instead of “scroll down”). The remaining 16.67% pertain to complex operations like “log in” or “register” that the LLM struggles to parse, some of which can be resolved by the exploration in the subsequent stage. Additionally, ‘Content fabrication’ represents the second largest category of errors, which is a notable issue associated with the LLM. It occasionally does not strictly adhere to the given step description and instead invents or elaborates on instructions.
@@ 5.4.3. Operation Mapping
The system achieves an 81.66% automatic prediction accuracy. As detailed in Figure 12, errors can be categorized as follows:
Data Loading Failures: 8.24% are caused by issues in page data loading, preventing the system from correctly identifying and operating widgets.
Semantic Ambiguity: 17.65% of errors arise from the system misidentifying widgets due to semantic ambiguities and fuzzy matching challenges. In this case, the agent can iterate attempts and refine solutions based on triggered results.
Agent-Related Errors: 18.82% of errors are due to the Grounding Agent taking an unrecoverable erroneous direction in task execution. 14.12% of errors happen when the Assessment Agent wrongly rejects operations initially predicted by the Grounding Agent. Such errors occur when contextual information is insufficient, but they diminish as the mobile interaction graph expands. The graph can provide the agents with a richer context for each operation’s implications.
Information Inadequacy: Errors due to incomplete page data constitute the largest proportion, accounting for 41.18%. Widgets without textual descriptions are particularly affected. To address this, the Mobile Semantics Agent employs advanced image-to-text models, although their ability to distinguish finer details among mobile widgets is limited. Fortunately, users can often provide context-specific interpretations to mitigate these issues. For example, they might interpret the icons, or clarify that “select” in a certain context means “switch” rather than “click”.

### v2

@@ 2. Clear cache
QQ Message 2 3~8 1. Share QQ screen
@@ 2. Provide feedback on an issue
TikTok Social Media 3 3~10 1. Cancel the monthly payment
@@ 2. View TikTok shop
Alipay Payment 4 3~8 1. Set the notification sound
@@ 2. View Alipay points
Weibo Social Media 5 4~7 1. Check creation earnings
@@ 2. View corrected followers
Lark Office 173 5~10 1. Log out
2. Add Lark reminders to the workspace
Beautiful Weather 366 3~7 1. Update the weather every 3 hours
Weather 2. Observe the weather in Beijing
Safe and Lifestyle 392 3~8 1. Set the shipping address
Sound 2. Enable the night mode
Film Entertainment 671 4~10 1. Join the user experience program
Encyclopedia 2. Bind QQ with the Film Encyclopedia
System Tools System - 4~8 1. Turn off the pure mode
@@ 5.2. Experiment Method
In our experiment, we evaluated three configurations of Prompt2Task:
  • Standard Prompt2Task.
  • Prompt2Task(AK): Prompt2Task enhanced with accumulated knowledge.
  • Prompt2Task(AK+minimal UI): Prompt2Task enhanced with accumulated knowledge and minimal necessary user intervention. Only essential user interventions during the operation mapping stage to provide important parameters are permitted. For instance, “Click a ringtone you like” in Figure 1 does require user intervention, while the scenario in Figure 10 does not inherently require it but arises due to the agent’s limitations.
  • Prompt2Task(AK+UI): Prompt2Task further enhanced with both accumulated knowledge and user intervention, providing perfect user intervention to maximize success rate.
The experiment assessed Prompt2Task’s performance at individual stages and across the entire pipeline. Notably, in a single stage where a task was not fully completed, without user intervention, the knowledge that Prompt2Task could accumulate was quite limited. Therefore, Prompt2Task(AK) was evaluated solely on its entire pipeline performance. When evaluating the performance of individual stages, we used ground-truth inputs, ensuring that any output deviation stemmed exclusively from the current stage, eliminating errors from preceding stages. Specifically, in the instruction generation stage, we input 100 ground-truth function and step descriptions; for the operation mapping stage, we input 100 ground-truth instruction lists. In evaluating the information collection stage and the entire pipeline, the input comprised 2,500 original textual prompts. To minimize the impact of input order and the inherent variability of LLM outputs, we randomly shuffled the inputs and conducted five rounds of testing to calculate the average performance.
We selected appropriate baselines for comparison at both individual stages and the entire process. For the information collection stage, manual queries on Google, Bing, and Baidu served as the baseline. For instruction generation, we used Seq2act (Li et al., 2020a) as the baseline. For the operation mapping stage, we employed strict text matching as the baseline. In assessing the entire pipeline, our baseline was using GPT-4 to generate instructions, followed by strict text matching for operation mapping. Additionally, considering some works that use step descriptions as input, we used LLM4Mobile (Wang et al., 2023) and AdbGPT (Feng and Chen, 2024) as baselines for the last two stages combined.
@@ 5.3. Experiment Results
The experiment results are organized based on the three stages within Prompt2Task as well as the performance of the entire pipeline.
@@ 5.3.1. Information Collection
This stage focuses on evaluating whether Prompt2Task can successfully generate a function description and step description based on a user-provided textual prompt.
Among the 2,500 textual prompts, 100 contained complete step information and didn’t require further online retrieval. Users accepted the analyzed results for 97.00% of these prompts, with a few exceptions due to the LLM’s data fabrication issues. The remaining 2,400 prompts were used to assess the top 5 tutorials retrieved by the system. The criteria for a “hit” was whether users agreed that a tutorial matched the ground-truth function description, regardless of on-phone executability. A tutorial at the top indicated that the system could automatically select it, eliminating the need for user intervention. Among the 2,400 prompts, 12.21% have incomplete descriptions of functions and required further supplementation.
@@
Table 4. Performance of various search engines in retrieving tutorials, using ground-truth function descriptions as input.
@@
Table 5. Performance of information collection, using 2,400 textual prompts as input.
@@
Manual(Baidu) 97.21 1.42 (sd=0.76) 87.96 -
Prompt2Task 98.83 1.27 (sd=0.91) 89.38 88.38
Prompt2Task(AK+UI) 100.00 1.22 (sd=0.74) 90.83 99.21
However, prompts provided by users often lack accuracy, leading to a decrease in effectiveness when using the original prompts for manual searches on these websites. Using them as baselines, we tested the performance of Prompt2Task. As shown in Table 5, Prompt2Task improved the Top-5 hit rate from 97.21% (Baidu) to 100%, while achieving a Top-1 hit rate of 90.83%. On average, each prompt required 0.15 user interventions (sd = 0.52), which represents the number of dialogue turns between the user and Prompt2Task (excluding the initial input). Remarkably, without user interventions, Prompt2Task demonstrated a Top-1 hit rate of 89.38%, surpassing the baseline performance.
After the agent matches function descriptions with online tutorials, it extracts concise step descriptions from complex web data. The accuracy of step descriptions refers to the correct extraction from web pages, ensuring they are free from fabrication, omissions, or errors. When Prompt2Task automatically selected the first candidate, the accuracy was 88.38%. When users could select from five candidates, the accuracy rose to 99.21%.
@@ 5.3.2. Instruction Generation
This stage evaluates Prompt2Task’s ability to transform step descriptions into a list of instructions in the predefined format.
Using the ground-truth step descriptions of the selected 100 Prompt2Task models as input, we evaluated correctness using two metrics, focusing solely on textual interpretation:
  • Complete Accuracy: The percentage of Prompt2Task models whose all generated instructions match the ground truth.
We used the Seq2Act model (Li et al., 2020a), a transformer trained using the ANDROIDHOWTO dataset, as our baseline (after translating step descriptions into English). Table 6 reveals that Prompt2Task achieved 100% accuracy for both complete and partial metrics. On average, achieving perfection required 0.93 user interventions per task (sd=1.34). It’s worth noting that while perfection necessitated user interventions, the subsequent pipeline stages are designed to accommodate minor errors, reducing the need for strict user oversight. Without user intervention, Prompt2Task attained 84.54% partial and 73.00% complete accuracy. In cold-start scenarios, Prompt2Task has achieved performance comparable to that achieved by related works (Li et al., 2020a; Vu et al., 2023), which require substantial training data (at least 10K-level).
@@
Partial Accuracy(%) Complete Accuracy(%) User Intervention
Seq2act (Li et al., 2020a) 70.13 54.00 -
Prompt2Task 84.54 73.00 -
Prompt2Task(AK+UI) 100.00 100.00 0.93 (sd = 1.34)
@@ 5.3.3. Operation Mapping
This stage evaluates Prompt2Task’s ability to execute operations on smartphones, guided by instructions. We used the ground-truth instruction lists of the 100 Prompt2Task models as input.
Our baseline for comparison utilized strict text matching to identify operation widgets. Table 7 shows the results. With user intervention (i.e., users can demonstrate any operations), the success rate reached 96.00%. The remaining 4% of tasks were unfeasible under the current test environment. On average, each task required 0.68 user interventions (sd = 0.85). Even without user assistance, the success rate was 62.00%, far surpassing the baseline. When users were allowed to provide necessary parameters, the success rate was 91.00%, indicating that 29.00% of the tasks require user-provided information even under optimal conditions.
In terms of the detailed performance without user intervention, Table 8 presents the frequency and accuracy of the Grounding Agent’s strategies when handling new instructions. The agent seldom uses the “expand” strategy, which also shows the lowest accuracy. This indicates that even though some instructions are wrong or outdated, they still hold great reference value. In most cases, the agent prefers to use “redirect” or “add” strategies to adapt to the current operating environment. Table 9 shows the frequency and accuracy of the Assessment Agent’s strategies for the operations predicted by the Grounding Agent. The agent lacked confidence in 16.53% of the decisions, of which 68.57% required necessary user intervention to complete the task. In the remaining cases, while user intervention is optional, it can help the agent make more informed decisions.
Table 7. Performance of operation mapping. Prediction accuracy is measured as the proportion of correctly executed operations out of the total number of operations, where an operation is considered correct if it is appropriate in the context at that time. The success rate is determined by the ratio of completed tasks.
Prediction Success User
Accuracy(%) Rate(%) Intervention
Text only 61.49 44.00 -
Prompt2Task 80.25 62.00 -
Prompt2Task(AK+minimal UI) 88.06 91.00 0.61 (sd = 0.77)
Prompt2Task(AK+UI) 92.11 96.00 0.68 (sd = 0.85)
Table 8. Frequency and accuracy of the Grounding Agent’s strategies. The “block” strategy indicates the agent could not predict the operation, thus it has no accuracy measure.
Frequency(%) Accuracy(%)
Strict Text Matching 26.44 100.00
Redirect 30.17 90.72
Add 39.50 56.69
Skip 1.24 100.00
Expand 1.40 33.33
Block 1.24 -
Table 9. Frequency and accuracy of the Assessment Agent’s strategies.
Frequency(%) Accuracy(%)
Follow 81.89 83.46
Change 5.04 53.13
Retract 3.31 95.24
Finish 9.76 100.00
@@ 5.3.4. The Entire Pipeline
To evaluate the overall effectiveness of Prompt2Task, we input 2,500 textual prompts into the pipeline, which comprises information collection, instruction generation, and operation mapping.
Given the task-related textual inputs, we did not use UGIF (Venkatesh et al., 2022) and CogAgent (Hong et al., 2023a) as baselines, as it would be unreasonable to require them to support our tasks without pre-provided tutorial sets or fine-tuning data. Additionally, we did not use Voicify (Vu et al., 2023), as our verification showed that only 15% of the tasks’ target pages had intent filters configured in the app code, which means that it is impossible to reach the remaining task pages using only page descriptions. Therefore, we established our baseline using GPT-4 for direct instruction generation from user inputs, coupled with strict text matching for operations.
As shown in Table 10, the success rate for each stage is strictly determined by the complete accuracy achieved up to that stage. Importantly, the overall success rate is not a simple product of success rates at each stage, due to the system’s built-in flexibility for error correction in downstream stages. For instance, the operation mapping stage doesn’t strictly follow the generated instructions; it also utilizes the agent’s exploratory strategies. For Prompt2Task(AK+minimal UI), the average number of user interventions per task is 0.42 (sd = 1.32), excluding the initial user input. For Prompt2Task(AK+UI), the average number is 0.69 (sd = 1.31). Notably, this number is smaller than the sum of user interventions required for each stage as reported earlier, which shows that the accumulation of knowledge from completed tasks can reduce user interventions.
Additionally, we compared Prompt2Task with baselines that support only step description inputs. We tested the performance using ground-truth step descriptions as inputs, which eliminates errors in online tutorial retrieval. Notably, it does not guarantee that the ground-truth tutorial is entirely feasible in the current test environment. While 31% of the tasks involve outdated step descriptions, only 11.82% of the individual steps are outdated, which means that a significant portion of the steps are still correct, and only minor corrections may be needed to complete the tasks. As shown in Table 11, Prompt2Task handles outdated tutorials effectively, achieving a 96% success rate when user intervention is applied. The discussion on outdated tutorials is detailed in Section 7.1. Moreover, the prediction accuracy of Prompt2Task(AK+minimal UI) and Prompt2Task(AK+UI) is lower than the success rate because the system executed redundant exploratory steps, which also led to a decrease in relative efficiency. However, for the goal of fully completing the task, this trade-off is considered acceptable.
Table 10. Success rate (%) across various stages, using 2,500 textual prompts as input.
@@
Prompt2Task 88.42 64.92 60.80
Prompt2Task(AK) 92.26 72.32 78.16
Prompt2Task(AK+minimal UI) 93.16 73.36 90.68
Prompt2Task(AK+UI) 99.12 99.12 95.24
Table 11. The performance of the entire system with ground-truth step descriptions as input. The definitions of prediction accuracy and success rate are the same as those in Table 7. Relative efficiency is the ratio of ground-truth operations to actual operations performed.
Prediction Accuracy(%) Success Rate(%) Relative Efficiency(%)
LLM4Mobile (Wang et al., 2023) 54.58 23.00 89.68
AdbGPT (Feng and Chen, 2024) 65.97 36.00 87.86
Prompt2Task 80.15 62.00 85.03
Prompt2Task(AK+minimal UI) 87.78 91.00 76.23
Prompt2Task(AK+UI) 91.46 96.00 76.13
@@ 5.4. Error Analysis
We analyzed the errors that occurred in the three stages of Prompt2Task.
@@ 5.4.1. Information Collection
Prompt2Task effectively matches tutorial web pages but falls short in content extraction. Two main issues affect the performance. First, web pages containing excessively long text compromise the LLM’s extraction capabilities, leading to information loss and increased risk of generating inaccurate content. Second, the LLM also tends to generate fabricated step descriptions when interpreting the textual prompts.
To address these challenges, we have considered two options for improving text quality. Initially, we considered implementing a specialized assessment agent, similar to the one used in the operation mapping, to manage text quality. However, we ultimately chose a human-in-the-loop approach, considering the efficiency of manual text assessment and the extra time needed for an agent-based solution. Our current system allows users to manually select high-quality step descriptions for subsequent stages.
@@ 5.4.2. Instruction Generation
Prompt2Task’s automatic instruction generation achieves a partial accuracy of 84.54%, with errors distributed across six main categories as shown in Figure 13. These errors are generally resolved through simple user corrections. Specifically, 43.93% of errors are due to instructions that go beyond the existing definition (Table 4.1.3). Within this subset, 24.51% involve the LLM using near-synonyms (e.g., “drag down” instead of “scroll down”). The remaining 19.42% pertain to complex operations like “log in” or “register” that the LLM struggles to parse, which can be potentially resolved by the exploration in the subsequent stage. Additionally, “Content fabrication” represents the second largest category of errors, which is a notable issue of the LLM. It occasionally does not strictly adhere to the given step description and instead invents or elaborates on instructions.
@@
Figure 12. Errors of instruction generation.
Figure 13. Errors of operation mapping.
@@ 5.4.3. Operation Mapping
Prompt2Task achieves an 88.06% prediction accuracy with minimal necessary user interventions. As detailed in Figure 13, errors can be categorized as follows:
Data Loading Failures: 9.89% are caused by issues in page data loading, making the system fail to access GUI widgets.
Semantic Ambiguity: 15.38% of errors arise from the system misidentifying widgets due to semantic ambiguities and fuzzy matching challenges. In this case, the agent can iterate attempts and refine the operation sequence based on triggered results.
Agent-Related Errors: 17.58% of errors are due to the Grounding Agent taking an unrecoverable erroneous direction in task execution. 13.19% of errors happen when the Assessment Agent wrongly rejects operations initially predicted by the Grounding Agent. Such errors occur when contextual information is insufficient, but they diminish as the mobile interaction graph expands. The graph can provide the agents with a richer context for each operation’s implications.
Information Inadequacy: Errors due to incomplete information constitute the largest proportion, accounting for 43.96%. Widgets without textual descriptions are particularly affected. Although the mobile semantics module in the Grounding Agent employs advanced image-to-text models, their ability to recognize the fine-grained semantics of widgets is limited. Fortunately, users can help interpret the widgets by demonstrating how the instructions map to the corresponding widgets. User demonstration can also provide context-specific interpretations, such as clarifying that “select” in a certain context means “switch” rather than “click”.